Kevin is 63 years old and Stephanie is 9 years old. How many years will it take until Kevin is only 4 times as old as Stephanie?
Answer by ewatrrr(24785)

Kevin is 63 years old and Stephanie is 9 years old.
x being the years i6 will take
63 + x = 4(9 + x)
27 = 3x
9 = x, years it will take
And... 9 years from NOW.
72 = 4*18